OpenStudy (trisaba):

make a polynomial where 1. the leading coefficient is one 2. the others are real and rational 3. the zeros are 3i and 2-i

OpenStudy (anonymous):

ok ready?

OpenStudy (trisaba):

yes

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

one zero is \(3i\) that makes another zero its conjugate \(-3i\) so the quadratic will be \[(x-3i)(x+3i)=x^2+9\]

OpenStudy (anonymous):

another way we could have found \(x^2+9\) is to set \[x=3i\] square both sides and get \[x^2=-9\] then add 9 to get \[x^2+9\] that takes care of the \(3i\) part

OpenStudy (trisaba):

i forgot about congegents lol

OpenStudy (anonymous):

we still have to deal with the \(2+i\) part

OpenStudy (anonymous):

there is a hard way to find the quadratic with zeros at \(2\pm i\), an easy way, and a real real easy way you pick

OpenStudy (trisaba):

r r easy

OpenStudy (anonymous):

lets do it the easy way first the real real easy way requires memorizing something (as with most math, the more you know the easier things are)

OpenStudy (anonymous):

set \[x=2+i\] and work backwards subtract 2 get \[x-2=i\] then square both sides (carefully) get \[(x-2)^2=i^2\] or \[x^2-4x+4=-1\]

OpenStudy (anonymous):

add 1 and get \[x^2-4x+5=0\] so your quadratic is \[x^2-4x+5\]

OpenStudy (anonymous):

the real real easy way is to know that if the solution is \(a+bi\) then the quadratic is \[x^2-2ax+a^2+b^2\] in your case \(a=2,b=1\) quadratic is \[x^2-2\times 2x+2^2+1^2=x^2-4x+5\]

OpenStudy (anonymous):

you still have one more job, to multiply \[(x^2+9)(x^2-4x+5)\] i will leave that to you

OpenStudy (anonymous):

hope it was more or less clear, if you have a question or two let me know

OpenStudy (trisaba):

thanks

OpenStudy (anonymous):

yw